# A dimension reduction approach to edge weight estimation for use in spatial models

### v1

@@ 5 Discussion
Within this article we introduced a framework for estimating edge weight matrices using the eigenvectors of the line graph’s Laplacian matrix. We show that this method may be used to enhance existing covariance models for areal data. We compared and contrasted the classically-used CAR model with the GDEF model introduced by Christensen and Hoff (2024), illustrating how both may be improved by this new framework. We evaluated properties of this method in a series of simulations and data examples. In general we find that the GDEF model is more flexible and capable of representing covariance structures with properties that are generally preferred by researchers, and that the introduced framework significantly improves the computational efficiency of the GDEF approach. We conclude this article with a brief discussion of potential avenues for future work and model extensions.
We are currently working on the development of an R package which may be used to easily implement our methodology. We also wish to evaluate how the INLA computing framework (Rue et al., 2009) may be used to complement or improve upon the model fitting approaches utilized within this article.
@@
Part of our assessment that the GDEF model exhibits more desirable behavior than the CAR model stems from findings that the latter characterizes fundamentally non-smooth spatial processes. Rue and Held (2005) and Lindgren et al. (2011) demonstrate that it is possible to define Markov random fields (MRFs) that approximate smooth spatial processes by allowing some edge weights to be negative. The space of non-CAR MRFs is worth exploring due to their computational advantages. It may be possible to adapt our method for edge weight estimation to allow for the inclusion of negative weights in sparse precision matrix formulations, though additional work would be necessary to ensure that a valid covariance matrix is always produced.
@@
While we have discussed estimation of W within the context of explicitly defining a covariance matrix, there are other settings in which a coherent framework for flexibly defining an edge weights matrix may be valuable. For example, resistance distance is a graph metric constructed using a matrix of non-negative edge weights that has seen widespread use in many statistical and ecological applications (Klein and Randić, 1993; Hanks and Hooten, 2013; Dickson et al., 2019). Even if not utilized for the purpose of defining distances or covariances between locations, an edge weights matrix may be viewed as characterizing the connectivity of different regions or entities within a network.
@@
An additional topic we would like to explore is how our method could be used to characterize the covariance of spatial processes on manifolds or restricted spatial domains. While most spatial models were been developed assuming that data are observed on some well-defined subset of $\mathbb{R}^{2}$, there is increased interest in the development of methods that can be applied to more complex domains in which Euclidean distances between observations do not accurately represent proximity with respect to the domain’s underlying geometry (Yang and Dunson, 2016). Mathematical distinctions exist between settings in which data are observed in Euclidean space but concentrated near a lower-dimensional manifold such as the surface of a sphere and settings in which which an $n$-dimensional spatial process exists only on an irregular subset of $\mathbb{R}^{n}$ due to underlying physical constraints (Dunson et al., 2022). Sangalli et al. (2013) provides an example of the latter setting in which population density on the Island of Montréal is modeled using a spatial process which is undefined beyond island borders and within the boundaries of an airport and industrial park. Whether data is observed on a true manifold or a restricted domain, we could define a graph based on the set of observed locations, defining edges between locations based on neighbors within a small radius or through use of the Delauney triangulation (Hjelle and Dæhlen, 2006). The LGL eigenvectors of this graph could then be utilized as in this article to define a valid covariance model over the complex spatial domain; furthermore, the flexibility of our edge weight estimation approach would ultimately allow covariance to be nonstationary with respect to the domain’s underlying geometry. Some model extensions would be required in order to allow covariance to be defined continuously between observations, as our present framework is defined only at the nodes of a graph. Insight from Lindgren et al. (2011) and Lee and Haran (2022) provide potential frameworks for solving this issue.

### v2

@@ 5 Discussion
Within this article we introduced a framework for estimating edge weight matrices using the eigenvectors of the line graph’s Laplacian matrix. Utilizing matrix calculations for the derivatives of the log likelihood with respect to basis function coefficients, we show that this method may be fit efficiently and how it can be used to enhance existing covariance models for areal data. We also evaluated properties of our approach in a series of simulations and examples that demonstrated its capacity to represent complex covariance structures while significantly improving the computational efficiency of the GDEF model introduced by Christensen and Hoff, (2024). We conclude this article with a brief discussion of potential avenues for future work and model extensions.
@@
Part of our preference for using the GDEF model in conjunction with the LGL basis rather than the CAR model stems from findings that the latter characterizes fundamentally non-smooth spatial processes. Rue and Held, (2005) and Lindgren et al., (2011) demonstrate that it is possible to define Markov random fields (MRFs) that approximate smooth spatial processes by allowing some edge weights to be negative. The space of non-CAR MRFs is worth exploring due to their computational advantages. It may be possible to adapt our method for edge weight estimation to allow for the inclusion of negative weights in sparse precision matrix formulations, though additional work would be necessary to ensure that a valid covariance matrix is always produced.
@@
While we have discussed estimation of W within the context of explicitly defining a covariance matrix, there are other settings in which a coherent framework for flexibly defining an edge weights matrix may be useful. For example, resistance distance is a graph metric constructed using a matrix of non-negative edge weights that has seen widespread use in many statistical and ecological applications (Klein and Randić,, 1993; Hanks and Hooten,, 2013; Dickson et al.,, 2019). Outside the context of distances and covariance, an edge weights matrix can be used to characterize the connectivity of different entities within any network.
@@
An additional topic we would like to explore is how our method could be used to characterize the covariance of spatial processes on manifolds or restricted spatial domains. While most spatial models have been developed assuming that data are observed on some well-defined subset of $\mathbb{R}^{2}$, there is increased interest in the development of methods that can be applied to more complex domains in which Euclidean distances between observations do not accurately represent proximity with respect to the domain’s underlying geometry (Yang and Dunson,, 2016). Mathematical distinctions exist between settings in which data are observed in Euclidean space but concentrated near a lower-dimensional manifold such as the surface of a sphere and settings in which an $n$-dimensional spatial process exists only on an irregular subset of $\mathbb{R}^{n}$ due to underlying physical constraints (Dunson et al.,, 2022). Whether data is observed on a true manifold or a restricted domain, we could define a graph based on the set of observed locations, defining edges between locations based on neighbors within a small radius or through use of the Delauney triangulation (Hjelle and Dæhlen,, 2006). The LGL eigenvectors of this graph could then be utilized as in this article to define a valid covariance model over the complex spatial domain; furthermore, the flexibility of our edge weight estimation approach would ultimately allow covariance to be nonstationary with respect to the domain’s underlying geometry.
Supplemental Materials
File containing additional simulation results, details on computation, and further discussion of relevant topics. Supplemental materials available online.